Please help me solve: a football is kicked vertically upward from a height of 2 feet with an initial speed of 65 feet per second. the formula h=2+65t-16t^2 describes the balls height above the ground h, in feet, t seconds after it was kicked. use this formula to find the balls height 1 second after it was kicked
Answer by ikleyn(52798)

.
Substitute t= 1 into the formula. You will get


    h = 2 + 65*1 - 16*1^2 = 2 + 65 - 16.


Now calculate it.


Happy calculations !